Business and finance. Paul invested $5000 in a time deposit. What interest will he earn for 1 year if the interest rate is 6.5%? Identify the rate, base, and amount in this problem.
Answer by chitra(359)

It is given that Paul invested $5000 for one year at the rate of 6.5%

So here,

Amount = $5000

Rate = 6.5 % = 0.065

Time = 1 yr

So, to calculate the interest, we use the formula for Simple Interest(SI).

SI = (P * T * R)/100

SI = (5000 * 1 * 6.5)/ 100

SI = (5000 * 1 * 0.065)

SI = 325.0

==> SI = $ 325


This shows that he earned an interest of $ 325 in a span of one year for the amount $5000 he deposited at the rate of 6.5 %

Hence, the solution.

Regards